Can you help me with this problem? Add in the lowest terms:

13/18 + 11/18 - Explain your answer in detail if you can.
Thanks!

Answer by KneesAndToes(13)

I'm not sure what you mean by add in lowest terms, but I saw this question was reposted, so I want to try and help. I'm going to guess that it should just be the answer given in lowest terms, because the problem is already in lowest terms:


13/18 + 11/18 = ? ... (add numerators)


13/18 + 11/18 = 24/18


Now you will need to put this answer in lowest terms. Unless common factors are obvious, I find it easiest to lower the numbers in steps. In this problem, both the numerator and the denominator are divisible by two:


24/18 = 12/9


These numbers are both divisible by three:


4/3 - This is the answer in lowest terms.


I found this website... it looks like it may be helpful to you:


http://www.helpwithfractions.com/index.html